# PSZ: Enhancing the SZ Scientific Lossy Compressor With Progressive Data Retrieval

### v1

@@ Abstract.
Compression is a crucial solution for data reduction in modern scientific applications due to the exponential growth of data from simulations, experiments, and observations. Compression with progressive retrieval capability allows users to access coarse approximations of data quickly and then incrementally refine these approximations to higher fidelity. Existing progressive compression solutions suffer from low reduction ratios or high operation costs, effectively undermining the approach’s benefits. In this paper, we propose the first-ever interpolation-based progressive lossy compression solution that has both high reduction ratios and low operation costs. The interpolation-based algorithm has been verified as one of the best for scientific data reduction, but previously no effort exists to make it support progressive retrieval. Our contributions are three-fold: (1) We thoroughly analyze the error characteristics of the interpolation algorithm and propose our solution IPComp with multi-level bitplane and predictive coding. (2) We derive optimized strategies toward minimum data retrieval under different fidelity levels indicated by users through error bounds and bitrates. (3) We evaluate the proposed solution using six real-world datasets from four diverse domains. Experimental results demonstrate our solution archives up to $487\%$ higher compression ratios and $698\%$ faster speed than other state-of-the-art progressive compressors, and reduces the data volume for retrieval by up to $83\%$ compared to baselines under the same error bound, and reduces the error by up to $99\%$ under the same bitrate.
@@ 1. Introduction
The increasing volume of scientific data generated by simulations, instruments, and observations has outpaced the data processing, storage, and transfer capabilities of modern computer systems including both workstations and supercomputers. For example, in climate research, the Coupled Model Intercomparison Project (CMIP) (Coupled Model Intercomparison Project, 2024) aims to advance our understanding of the climate system by coordinating standardized experiments with Earth System Models (ESMs) and thus enabling comprehensive comparisons of how different models represent past, present, and future climate conditions. Thanks to the rapid evolution of leading HPC systems, each successive phase of CMIP has seen significant increases in data volume – CMIP3 generated around 40 TB (Abdulah et al., 2024; Balaji et al., 2018), CMIP5 about 2 PB (Balaji et al., 2018), and CMIP6 (Coupled Model Intercomparison Project Phase 6 dashboard, 2024) exceeding 28 PB. Such an increase in volume poses unprecedented challenges to store, process, and analyze the data.
@@
Motivation: Although many scientific lossy compression solutions have been proposed in the recent decade, most of them only support decompression to a single fidelity level once the data is compressed, which restricts their broad utilization. On one hand, scientific analyses often require different data fidelity levels as the subject of the study or the stage of investigation has diverse tolerance for data precision. For example, in hydrodynamic simulation, reconstructing viscosity may require a $2^{-5}$ finer precision compared with reconstructing vorticity from the same data field (Magri and Lindstrom, 2024). As a result, without progressive capability, researchers must compress data conservatively at the highest possible fidelity – even though only a handful of analyses truly require it – ultimately diminishing the overall effectiveness of data reduction. On the other hand, when analyzing multiple snapshots, fields, or regions, researchers often first identify patterns or areas of interest at a coarse level before committing resources to detailed analysis (Magri and Lindstrom, 2024; Liang et al., 2021; Wu et al., 2024). Without the ability of progressive retrieval, scientists have to always load the entire compressed data and decompress it at full precision. This not only increases the time and resources required for data loading and decompression but also delays subsequent analyses and scientific discoveries.
@@
Limitation of state-of-art approaches: Despite the necessity of progressive retrieval in scientific compression as highlighted in above scenarios, supporting this functionality is challenging due to several factors. First, achieving both high fidelity and high compression ratios simultaneously is challenging, as these objectives often depend on fundamentally different algorithms, where optimizing one may come at the expense of the other. Second, straightforward progressive techniques often introduce significant operational overhead, requiring multiple passes of decompression and reprocessing – contradicting the goal of progressive decompression which is to save time and resources. Third, it is already non-trivial to guarantee error bounds for partial decompression, not alone to say the progressive technique which requires the data accumulated from multiple levels to be within acceptable error margins. Consequentially, few such progressive solutions exist, and they often fall short due to low compression ratios, high operational costs, and a lack of stringent error restrictions. These limitations hinder their adoption in scientific applications.
@@
Key insights and contributions: In this paper, we propose the first high ratio, fast, and error-bounded progressive lossy compression solution, based on the interpolation algorithm. Our contributions are three-fold:
@@
  • We thoroughly analyze the error characteristics to build a progressive compressor based on the interpolation algorithm, and propose our progressive solution, IPComp, based on the prediction model with multi-level bitplane and predictive coding. Besides having high effectiveness on retrieval, our solution supports retrieval under arbitrary error-bound settings, and it only executes decompression once for each retrieval request, compared with residual-based alternatives, which support a limited number of error bounds and require multiple passes of decompression for a single request.
@@
  • Experimental methodology and artifact availability: We evaluate the proposed solution using six real-world datasets from four diverse domains over four state-of-the-art compressors. Experimental results demonstrate our solution archives xx higher compression ratios and up to $487\%$ higher compression ratios and $698\%$ faster speed than other state-of-the-art progressive compressors, and reduces the data volume for retrieval by up to $83\%$ compared to baselines under the same error bound, and reduces the error by up to $99\%$ under the same bitrate. The source code of IPComp will be available to the public upon acceptance of the paper.
@@ 3.1.2. Objectives
Symbol Description $n$ Numbers of elements of the input $x$ The input dataset $\hat{x}$ The decompressed output with lossy error $y,\hat{y}$ The decorrelated data and its lossy version q The quantized data e The encoded data (saved as compressed data) $eb$ The bound of maximum lossy error T, P The transform and prediction function Q The quantization function E The coding function rm $\|\cdot\|_{\infty}$ $L_{\infty}$ norm - the maximum absolute value of the input $V$ n-dimensional linear space $V_{l}$ The sublinear space of the l-th level. $V_{l}$ and $V_{m}$ are orthogonal to each other when $l\neq m$ $\Pi_{l}$ The $L_{2}$ norm projection operator of $V_{l}$ $x_{i},y_{i}$ The subscript $i$ denotes the i-th component of the vector. $x_{l}$ the vector is projected to $V_{l}$. $x_{l}=\Pi_{l}x$, which applies to any vector with subscription $l$
@@
  • High compression ratio: High compression ratio should be the fundamental goal for any compression solution. Some progressive designs, such as SZ3-M shown in Section 6, sacrifice compression ratio for progressiveness, but we argue that the low compression ratio will prevent the adoption of such solutions and thus make their progressive capability useless. Instead, we aim to reach compression ratios higher than or equal to all other SOTAs.
  • Progressive retrieval: Users should be able to retrieve an approximate version of $x$ at a lower fidelity level $F_{1}$ and incrementally refine it to higher fidelity levels $F_{2},\dots,F_{n}$ without the need for multiple decompression operations.
  • Fast speed: The speed of compression and decompression should be equal to or faster than all other SOTAs, and the decompression process should require only a single pass to retrieve data at any specified fidelity level $F_{i}$.
  • Error guarantee: Guarantee that for each fidelity level $F_{i}$, if $F_{i}$ has error bound restriction of $\epsilon_{i}$, the reconstruction error $E_{i}$ satisfies $E_{i}\leq\epsilon_{i}$. This ensures that the compressed data remains within acceptable accuracy limits for scientific computations.
To the best of our knowledge, there is no scientific compressor that simultaneously achieves all these objectives.
3.2. Overall Design
Figure 1. Overall design of our solution IPComp (the compressed data contains multiple independently decompressible blocks, represented as 1-5 in the diagram)
@@ 3.2. Overall Design
We introduce a progressive lossy compressor designed to efficiently meet the diverse precision and fidelity requirements of scientific applications. Our solution IPComp is based on the interpolation algorithm which has been verified as the leading non-progressive scientific reduction method, and we are the first to make it progressive. IPComp mainly has three innovative modules.
@@
We demonstrate the compression and decompression workflow of IPComp in Figure 1. For comparison, we also highlight the differences between our approach and SOTA residual-based methods.
The compression workflow begins with the original dataset, which is processed by an Interpolation Predictor to decorrelate the dataset based on spatial and numerical relationships, and quantize the output from floating point format to integers based on the error bound (Tao et al., 2017). Following this, the Predictive coder encodes the prediction integers by bitplanes into multiple independent groups, each contributing incrementally to the accuracy of the decompressed data.
The data retrieval workflow is designed to support multi-fidelity retrieval requests efficiently, leveraging an Optimized Data Loader that extracts only the required precision groups for a given query. Figure 1 demonstrates how the workflow works using three requests in increasing fidelity levels. Requests 1 asks for an error bound of 1E-2, so the data loader retrieves the first few precision groups (e.g., groups 1, 2) and passes them to the Interpolation Predictor and Progressive coder. These groups are processed together at one single pass to generate low-precision output data, suitable for quick exploration or coarse analyses. For the second request which indicates a targeting bitrate of 2, the data loader determines one additional precision group (e.g., group 3) is enough to satisfy such fidelity rule. As a result, the predictor and coder reconstruct this mid-precision representation of the dataset using group 3 on top of the low-precision result. Finally, the last request asks to retrieve all groups, so groups 4 and 5 are loaded consequently to build the high-precision reconstruction.
In contrast, residual-based SOTA approaches involve significant computational overhead and redundant operations. Those approaches will first compress in a large bound, then compress the residual from the last time using a smaller bound, and repeat such residual compression until a targeted bound is reached. As a result, they require multiple decompression passes to fulfill a single retrieval request. For example, in the SOTA workflow of Figure 1, the error bound for groups $1^{\prime}$, $2^{\prime}$, $3^{\prime}$ are 1, 1e-1, and 1e-2, respectively. For a single request of 1e-2, those approaches need to load three groups and execute the decompression three times to apply the result on top of each group.
@@
4. Progressive interpolation
In this section, we first analyze the interpolation-based algorithm, which is the leading non-progressive scientific compression strategy, and then we propose our solution to make it progressive by the prediction modal and predictive coder.
Figure 3. A typical lossy compression workflow (Berger and Gibson, 1998). T/P and Q represent decorrelation, quantization stages respectively. The quantization stage is lossy thus $\hat{y}$ is the lossy version of $y$. Definitions for $x$, $y$, etc can be found in Table 1
@@ 4.1. Introduction to none-progressive interpolation algorithm
Most scientific lossy compression workflows are composed of three key steps – decorrelation, quantization, and encoding (Zhao et al., 2021; Lindstrom, 2014; Li et al., 2023; Ainsworth et al., 2018). As shown in Figure 3, the input data is defined as a vector $x$ in a linear space $\mathbb{R}^{n}$, where $n$ represents the total elements in the data set. In the decorrelation step, a transform $T$ or prediction function $P$ is applied to $x$, resolving a decorrelated vector $y$. After a lossy quantization $Q$ on $y$, the quantized integers could be restored for further lossless compression. Accordingly, the decompression process executes the three steps reversely.
One critical aspect of designing an effective lossy compressor is to construct the best-suit decorrelation algorithm. Our solution IPComp uses interpolation-based decorrelation, which has been proven to be the leading solution in the scientific domain (Zhao et al., 2021; Liu et al., 2022, 2024; Ainsworth et al., 2018). Its core idea is to estimate unknown data points at fixed relative indices using interpolation formulas, such as linear interpolation and cubic spline interpolation. Unlike many other predictors (such as regression-base ones) that require storing coefficients for reconstruction, the interpolation approach eliminates such storage overhead by relying on predefined prediction formulas for fixed indices. This combination of high prediction accuracy and the elimination of coefficient storage enables interpolation-based methods to achieve much higher compression ratios than many alternative approaches (Zhao et al., 2021).
@@ 4.2.1. Transform models
Transform models (Berger and Gibson, 1998) can be viewed as a linear mapping between original data and decorrelate data. Many transform models, including Fourier transform, discrete cosine transform, wavelet transform, and orthogonal transform, have been adopted in lossy compression. For example, ZFP (Lindstrom, 2014) uses nearly orthogonal block transform, and MGARD (Ainsworth et al., 2018) is based on multilevel interpolation.
@@
The value of $L_{\infty}(T^{-1})$ depends on the specific transformation function. Taking the widely used transform function $x_{i}=x_{i}-x_{i-1}$ as an example, its corresponding transform and inverse transform formulas are:
$T=\begin{bmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&1\end{bmatrix},\quad T^{-1}=\begin{bmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 1&1&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&1&1&\cdots&1\end{bmatrix}$
In this case, the $L_{\infty}$ norm of $(T^{-1})$ equals the maximum sum of the rows in the matrix $L_{\infty}$, which is the total number of the elements $n$. As a result, the distortion can be bounded by n times the max error in the transformed domain.

### v2

@@ Abstract.
Compression is a crucial solution for data reduction in modern scientific applications due to the exponential growth of data from simulations, experiments, and observations. Compression with progressive retrieval capability allows users to quickly access coarse approximations of data and then incrementally refine these approximations to higher fidelity. Existing progressive compression solutions suffer from low reduction ratios or high operation costs, effectively undermining the approach’s benefits. In this paper, we propose the first-ever interpolation-based progressive lossy compression solution that has both high reduction ratios and low operation costs. The interpolation-based algorithm has been verified as one of the best for scientific data reduction, but previously, no effort exists to make it support progressive retrieval. Our contributions are three-fold: (1) We thoroughly analyze the error characteristics of the interpolation algorithm and propose our solution, IPComp, with multi-level bitplane and predictive coding. (2) We derive optimized strategies toward minimum data retrieval under different fidelity levels indicated by users through error bounds and bitrates. (3) We evaluate the proposed solution using six real-world datasets from four diverse domains. Experimental results demonstrate our solution archives up to $487\%$ higher compression ratios and $698\%$ faster speed than other state-of-the-art progressive compressors, and reduces the data volume for retrieval by up to $83\%$ compared to baselines under the same error bound, and reduces the error by up to $99\%$ under the same bitrate.
@@ 1. Introduction
The increasing volume of scientific data generated by simulations, instruments, and observations has outpaced the data processing, storage, and transfer capabilities of modern computer systems, including both workstations and supercomputers. For example, in climate research, the Coupled Model Intercomparison Project (CMIP) (Coupled Model Intercomparison Project, 2024) aims to advance our understanding of the climate system by coordinating standardized experiments with Earth System Models (ESMs) and thus enabling comprehensive comparisons of how different models represent past, present, and future climate conditions. Thanks to the rapid evolution of leading HPC systems, each successive phase of CMIP has seen significant increases in data volume – CMIP3 generated around 40 TB (Abdulah et al., 2024; Balaji et al., 2018), CMIP5 about 2 PB (Balaji et al., 2018), and CMIP6 (Coupled Model Intercomparison Project Phase 6 dashboard, 2024) exceeding 28 PB. Such an increase in volume poses unprecedented challenges to store, process, and analyze the data.
@@
Motivation: Although many scientific lossy compression solutions have been proposed in the recent decade, most of them only support decompression to a single fidelity level once the data is compressed, which restricts their broad utilization. On the one hand, scientific analyses often require different data fidelity levels as the subject of the study or the stage of investigation has diverse tolerance for data precision. For example, in hydrodynamic simulation, reconstructing viscosity may require a $2^{-5}$ finer precision compared with reconstructing vorticity from the same data field (Magri and Lindstrom, 2024). As a result, without progressive capability, researchers must compress data conservatively at the highest possible fidelity – even though only a handful of analyses truly require it – ultimately diminishing the overall effectiveness of data reduction. On the other hand, when analyzing multiple snapshots, fields, or regions, researchers often first identify patterns or areas of interest at a coarse level before committing resources to detailed analysis (Magri and Lindstrom, 2024; Liang et al., 2021; Wu et al., 2024). Without the ability of progressive retrieval, scientists have to always load the entire compressed data and decompress it at full precision. This not only increases the time and resources required for data loading and decompression but also delays subsequent analyses and scientific discoveries.
@@
Limitation of state-of-art approaches: Despite the necessity of progressive retrieval in scientific compression, as highlighted in the above scenarios, supporting this functionality is challenging due to several factors. First, achieving both high fidelity and high compression ratios simultaneously is challenging, as these objectives often depend on fundamentally different algorithms, where optimizing one may come at the expense of the other. Second, straightforward progressive techniques often introduce significant operational overhead, requiring multiple passes of decompression and reprocessing – contradicting the goal of progressive decompression, which is to save time and resources. Third, it is already non-trivial to guarantee error bounds for partial decompression, not alone to say the progressive technique which requires the data accumulated from multiple levels to be within acceptable error margins. Consequentially, few such progressive solutions exist, and they often fall short due to low compression ratios, high operational costs, and a lack of stringent error restrictions. These limitations hinder their adoption in scientific applications.
@@
Key insights and contributions: In this paper, we propose the first high ratio, fast, and error-bounded progressive lossy compression solution based on the interpolation algorithm. Our contributions are three-fold:
@@
  • We thoroughly analyze the error characteristics to build a progressive compressor based on the interpolation algorithm and propose our progressive solution, IPComp, based on the prediction model with multi-level bitplane and predictive coding. Besides having high effectiveness on retrieval, our solution supports retrieval under arbitrary error-bound settings, and it only executes decompression once for each retrieval request, compared with residual-based alternatives, which support a limited number of error bounds and require multiple passes of decompression for a single request.
@@
  • Experimental methodology and artifact availability: We evaluate the proposed solution using six real-world datasets from four diverse domains over four state-of-the-art compressors. Experimental results demonstrate our solution archives up to $487\%$ higher compression ratios and $698\%$ faster speed than other state-of-the-art progressive compressors and reduces the data volume for retrieval by up to $83\%$ compared to baselines under the same error bound and reduces the error by up to $99\%$ under the same bitrate. The source code of IPComp will be available to the public upon acceptance of the paper.
@@ 3.1.2. Objectives
Symbol Description $n$ Numbers of elements of the input $x$ The input dataset $\hat{x}$ The decompressed output with lossy error $y,\hat{y}$ The decorrelated data and its lossy version q The quantized data $eb$ The bound of maximum lossy error T, P The transform and prediction function Q The quantization function $\|\cdot\|_{\infty}$ $L_{\infty}$ norm - the maximum absolute value of the input $V_{l}$ The sublinear space of the l-th level. $V_{l}$ and $V_{m}$ are orthogonal to each other when $l\neq m$ $\Pi_{l}$ The $L_{2}$ norm projection operator of $V_{l}$ $x_{i},y_{i}$ The subscript $i$ denotes the i-th component of the vector. $x_{l}$ the vector is projected to $V_{l}$. $x_{l}=\Pi_{l}x$, which applies to any vector with subscription $l$
Figure 1. A typical lossy compression workflow. T/P represents decorrelation and Q means quantization. The quantization stage is lossy thus $\hat{y}$ is the lossy version of $y$. Definitions for $x$, $y$, etc can be found in Table 1
@@
  • High compression ratio: We aim to reach compression ratios higher than or equal to all other SOTAs, as a high compression ratio (CR) is essential for compressors to be effectively utilized in the scientific field. Some progressive designs, such as SZ3-M shown in Section 6, sacrifice compression ratio for progressiveness, but we argue that the low compression ratio will prevent the adoption of such solutions and thus make their progressive capability useless.
  • Progressive retrieval: We aim to support users to retrieve the output at a lower fidelity level $F_{1}$ and incrementally refine it to higher fidelity levels $F_{2},\dots,F_{n}$ without the need for multiple decompression passes.
  • Fast speed: Our speed of compression and decompression should be equal to or faster than all other SOTAs, and the decompression process should require only a single pass to retrieve data at any specified fidelity level $F_{i}$.
  • Error guarantee: We guarantee that for each fidelity level $F_{i}$, if $F_{i}$ has error bound restriction of $\epsilon_{i}$, the reconstruction error $E_{i}$ satisfies $E_{i}\leq\epsilon_{i}$. This ensures that the compressed data remains within acceptable accuracy limits for scientific computations.
To the best of our knowledge, no compressor simultaneously achieves all these objectives yet.
3.2. Overview of IPComp
Figure 2. Overall design of our solution IPComp (the compressed data contains multiple decompressible blocks, represented as 1-5 in the diagram)
@@ 3.2. Overview of IPComp
We introduce IPComp, a progressive lossy compressor designed to efficiently meet the diverse precision and fidelity requirements of scientific applications. IPComp is based on the interpolation algorithm which has been verified as the leading non-progressive reduction method, and we are the first to make it progressive. IPComp mainly has three innovative modules, shown in Figure 2.
@@
We demonstrate the compression and decompression workflow of IPComp in Figure 2. For comparison, we also highlight the differences between IPComp and residual-based SOTAs.
The compression workflow begins with the original dataset, which is processed by an Interpolation Predictor to decorrelate the dataset based on spatial and numerical relationships, and quantize the output from floating point format to integers based on the error bound. Following this, the Predictive coder encodes the prediction integers by bitplanes into multiple independent blocks, each contributing incrementally to the accuracy of the decompressed data.
The data retrieval workflow is designed to support multi-fidelity retrieval requests efficiently, leveraging an Optimized Data Loader that extracts only the required blocks for a given query. Figure 2 demonstrates how the workflow works using three requests in increasing fidelity orders. Requests 1 asks for an error bound of 1E-2, so the data loader retrieves the first few precision blocks (e.g., blocks 1, 2) and passes them to the Interpolation Predictor and Progressive coder. These blocks are processed together at one single pass to generate low-precision output, suitable for quick exploration or coarse analyses. For the second request which indicates a targeting bitrate of 2, the data loader determines one additional block (e.g., block 3) is optimal for such a request. As a result, the predictor and coder reconstruct this mid-precision representation using block 3 on top of the low-precision result. Finally, the last request asks to retrieve all blocks, so blocks 4 and 5 are loaded consequently to build the high-precision reconstruction.
In contrast, residual-based SOTA approaches involve significant computational overhead and redundant operations. Those approaches will first compress in a large bound, then compress the residual from the last time using a smaller bound, and repeat such residual compression until a targeted bound is reached. As a result, they require multiple decompression passes to fulfill a single retrieval request. For example, in the SOTA workflow of Figure 2, the error bound for blocks $1^{\prime}$, $2^{\prime}$, $3^{\prime}$ are 1, 1e-1, and 1e-2, respectively. For a single request of 1e-2, those approaches need to load three blocks and execute the decompression three times to apply the result on top of each block.
@@
4. Progressive design of IPComp
In this section, we first analyze the interpolation-based algorithm, which is the leading non-progressive scientific compression strategy, and then we propose our solution to make it progressive by the prediction model and predictive coder.
@@ 4.1. Introduction to none-progressive interpolation algorithm
Most scientific lossy compression workflows are composed of three key steps – decorrelation, quantization, and encoding (Zhao et al., 2021; Lindstrom, 2014; Li et al., 2023; Ainsworth et al., 2018). As shown in Figure 1, the input data is defined as a vector $x$ in a linear space $\mathbb{R}^{n}$, where $n$ represents the total elements in the dataset. In the decorrelation step, a transform $T$ or prediction function $P$ is applied to $x$, resolving a decorrelated vector $y$. After a lossy quantization $Q$ on $y$, the quantized integers are further lossless coded to bitstream as the final compressed data. Accordingly, the decompression process executes the three steps reversely.
One critical aspect of designing an effective lossy compressor is to construct the best-suit decorrelation algorithm. Our solution IPComp uses interpolation-based decorrelation, which has been proven to be the leading solution in the scientific domain (Zhao et al., 2021; Liu et al., 2022, 2024; Ainsworth et al., 2018). Its core idea is to estimate unknown data points at fixed relative indices using interpolation formulas, such as linear interpolation and cubic spline interpolation. Unlike many other predictors (such as regression-base ones) that store coefficients for reconstruction, the interpolation approach eliminates such storage overhead by relying on predefined prediction formulas for fixed indices. As a result, this approach achieves much higher compression ratios than many alternatives (Zhao et al., 2021).
@@ 4.2.1. Transform models
Transform models can be viewed as a linear mapping between original data and decorrelate data. Many transform models, including Fourier transform and wavelet transform, have been adopted in lossy compression. For example, ZFP (Lindstrom, 2014) uses nearly orthogonal block transform, and SPERR (Li et al., 2023) is based on the cdf97 wavelet.
@@
The value of $L_{\infty}(T^{-1})$ depends on the specific transformation function. Taking the basic but widely used transform function $x_{i}=x_{i}-x_{i-1}$ as an example, its corresponding transform and inverse transform formulas are:
$T=\begin{bmatrix}1&0&\cdots&0\\ -1&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\end{bmatrix},\quad T^{-1}=\begin{bmatrix}1&0&\cdots&0\\ 1&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\cdots&1\end{bmatrix}$
In this case, the $L_{\infty}$ norm of $(T^{-1})$ equals the maximum sum of the rows in the matrix $T$, which equals the input size $n$. As a result, any distortion in the transformed domain will be amplified in the original domain, and the maximum error could become $n$ times the maximum error in the transformed domain. Given that $n$ in most scientific datasets is at least $10^{7}$, such a degree of distortion amplification may be unacceptable for progressive compression.